Consider a perfectly competitive market in the short run, assume that the demand curve is given as shown below. P=100-4Q, also you are told that the total cost is given as TC=50+4Q+2Q^2 and the marginal cost MC=4+4Q. How many firms are in the industry in the short run?
1
Expert's answer
2019-06-11T10:47:00-0400

Firm supply=MC=4+4Q

 Equilibrium Price=52, Quantity=12

P=4+4q

52=4+4q, q=12

Each competitive firm produces 12 units hence there is one firm